OpenStudy (anonymous):

Solve x over 2 + 4 = 10

OpenStudy (anonymous):

\[x^2 + 4 = 10\] \[x^ = 10-4=6\] \[x = \pm \sqrt{6}\]

OpenStudy (anonymous):

Second line will be \[x^2 = 10-4=6\]

OpenStudy (anonymous):

thats not what she is asking @zarax

OpenStudy (anonymous):

she's asking:\[\frac{ x }{ 2 }+4=10\] @zarax

OpenStudy (anonymous):

so what you would do is multiply both sides by 2:\[x+8=20\]

OpenStudy (anonymous):

and then subtract 8 to get: x= 12